# The Smarr formula within the Geroch-Held-Penrose formalism

### v1

@@ IV Smarr relation in GHP formalism
Following Padmanabhan’s approach, we first study a simplified case taking $g(r)=f(r)^{-1}$. This choice of the metric function implies that $G^{t}_{t}=G^{r}_{r}$ (in fact, this is true for any static and spherically symmetric black hole horizon even if the metric is not determined by a single function [33]) and then the only two independent field equations are $G^{r}_{r}=T^{r}_{r}$ and $G^{\theta}_{\theta}=T^{\theta}_{\theta}$.
In terms of the Ricci coefficients, this choice of the metric implies that the only non-vanishing components are $\Phi_{11}$ and $\Lambda$. Thus, these coefficients must be related with the different components of the Einstein tensor. As we showed in Sec. III, the only nontrivial field equations are the ones given in the latter diagram, see Fig. 1. Interestingly,
@@
In this particular case, the spacetime is of type D and the congruences associated to the principal null directions are geodesic, non-rotating and shear-free. Then $\kappa=\omega=\sigma=0$, as is shown in App. A. This implies that $\rho$ is a real quantity that determines the congruence expansion of the principal null directions normal to the marginal surface [32]. At this point, Eq. (18) reads
@@
The first term in Eq. (24) is strongly related with the Hayward definition of quasi-local energy [19]
@@
where $A$ denotes the area of the black hole horizon and $\star$ denotes the Hodge operator defined on the horizon $H$. This coincides with the energy term obtained by Padmanabhan, $E=a/2$, for the static and spherically symmetric case [10].
Moreover, we give a generalized definition for the internal energy of a black hole following the Hayward’s procedure [34]. Over a compact 2-surface $S$, Hayward introduces the Hamiltonian 2-form
@@
where $A_{0}$ is a dimensionless constant representing the area of a unit horizon [^12] We use $\sqrt{A/A_{0}}$ as it is the most natural unit of length., being $A_{0}=4\pi|g-1|$ in the non-flat case (where $g$ is the genus of the horizon) and $A_{0}=L_{x}L_{y}$ in the flat torus case [^13] This is the unique compact surface of constant Gaussian curvature equal to zero., where $L_{x}$ and $L_{y}$ are the length of the coordinate ranges after the identifications $x\sim x+L_{x}$ and $y\sim y+L_{y}$. In particular, introducing the Euler characteristic of the horizon, denoted by $\chi(H)$, Eq. (27) can be written as $U=\frac{\chi(H)}{2}\left(\frac{a}{2}\right)$ which coincides with the Padmanabhan’s definition of internal energy [10].
The quantity $\mbox{}^{\prime}\rho$ is constant over a stationary horizon [18]. Moreover, by direct computation $\Phi_{11}^{ph}+3\Lambda^{ph}$ depends only on the $r$ coordinate and the Gaussian curvature of the marginal surface, which we assumed to be constant. Thus, after using Eq. (27) in Eq. (24), this can be rewritten as
@@
As we commented before, Hayward’s definition for trapping gravity [19], shown in Eq. (14) does not coincide with the surface gravity even in the context of static and spherically symmetric spacetimes , e.g. the Reissner-Nördstrom solution [20, 21]. Our definition in Eq. (31) overcomes this problem and is valid for any static spacetime foliated by compact orientable 2-surfaces with constant Gaussian curvature.
@@
At this point, some comments about the thermodynamic volume of the black hole should be made. We first analyze the vacuum case in GR, where $\Phi_{11}^{ph}=0$, which admits a cosmological constant. Such a case has been widely studied, identifying the presence of a cosmological constant with a $P\mathrm{d}V$ term in the first law of black hole thermodynamics [26]. Indeed, this allows the definition of the thermodynamic volume of the black hole, which may not coincide with the volume of a 3-manifold defined as having the horizon as its boundary (this 3-manifold may even be not well defined, for example in some hyperbolic cases). Thus, the thermodynamic volume should be defined as the conjugate thermodynamic variable to the pressure
@@
As commented in Sec. II, when a cosmological constant is considered, the mass of the black hole no longer coincides with its internal energy, but it is identified with the chemical enthalpy [26]. This is,
@@
However, when returning to the general case with $\Phi_{11}\neq 0$, the resulting Penrose-Rindler Smarr formula is again the Smarr relation, but the Riemann decomposition is not the physical decomposition as the mass of the black hole is not $M=U+PV$ when $P$ is defined as in Eq. (32). An explicit example is shown below.
@@
and noting that the left hand side term is equal to the black hole mass [35] this leads to
@@
For the vacuum case, the general action in $f(R)$ formalism is [36]
@@
Applying the variational principle, $\delta S=0$, the field equations for $f(R)$ gravity can be written as Eq. (8), where we have defined the effective stress-energy tensor as
with $F(R)\coloneqq df(R)/dR$.
@@
The first term coincides with the internal energy of the black hole in $f(R)$ gravity for the spherically symmetric case [37]
@@
Additionally, the temperature remains the same independently of the theory (this is a expected result as the temperature is entirely determined by the geometry of the horizon) after identifying the corresponding entropy [9]
@@
V Final remarks and comments
@@
$\dfrac{1}{4\pi}\sqrt{\dfrac{A}{A_{0}}}\int_{H}\star\left(k_{g}-(K+\overline{K})\right)=0\Leftrightarrow M=2TS+\Phi Q-2PV.$ (49)
@@
Even more, this framework allow us to study exotic topologies, obtaining some constraints on black hole existence in particular scenarios. For instance, the Penrose-Rindler Smarr formula implies that only spherical black holes with non-zero temperature can exist in General Relativity under the dominant energy condition. This recovers the well-known topology law for black holes [34]. Furthermore, the generalized definition of internal energy reveals that flat toroidal black holes have zero internal energy, while some hyperbolic black holes exhibit negative internal energy, which has been pointed out in previous works [35].
It is important to remark that the temperature we have proposed is not valid for the stationary axisymmetric case. In fact, despite the surface gravity over the Kerr black hole horizon must be constant due to the horizon topology, the quantity $\mbox{}^{\prime}\rho$ is not constant over the horizon (as it depends on angular coordinates), which suggests a revised generalization of trapping gravity for stationary spacetimes. The study of this procedure applied to a more general ansatz is deferred to future work, including the connection between the Penrose and Rindler formula and the Smarr formula in Kerr-Newmann black holes or even in more generic configurations.
@@ Acknowledgments
We acknowledge financial support from the Generalitat Valenciana through PROMETEO PROJECT CIPROM/2022/13. A. G. acknowledges Fundación Humanismo y Ciencia for financial support. A.G. acknowledges Martuan for continuous support.

### v2

@@ IV Smarr relation in GHP formalism
Following Padmanabhan’s approach, we first study a simplified case taking $g(r)=f(r)^{-1}$. This choice of the metric function implies that $G^{t}_{t}=G^{r}_{r}$ (in fact, this is true for any static and spherically symmetric black hole horizon even if the metric is not determined by a single function Medved2004 ) and then the only two independent field equations are $G^{r}_{r}=T^{r}_{r}$ and $G^{\theta}_{\theta}=T^{\theta}_{\theta}$.
In terms of the Ricci coefficients, this choice of the metric implies that the only non-vanishing components are $\Phi_{11}$ and $\Lambda$. Thus, these coefficients must be related with the different components of the Einstein tensor. As we showed in Sec. III, the only nontrivial field equations are the ones given in the latter diagram in Fig. 1. Interestingly,
@@
In this particular case, the spacetime is of type D and the congruences associated to the principal null directions are geodesic, shear-free and non-rotating. Then $\kappa=\sigma=0$ and $\omega=\mathrm{Im}(\rho)=0$, as is shown in App. A. This implies that $\rho$ is a real quantity that determines the congruence expansion of the principal null directions normal to the marginal surface GHPBook . At this point, Eq. (22) reads
@@
The first term in Eq. (28) is strongly related with the Hayward definition of quasi-local energy Hayward1994
@@
where $A$ denotes the area of the black hole horizon and $\star$ denotes the Hodge operator defined on the horizon $H$. This coincides with the energy term obtained by Padmanabhan, $E=a/2$, for the static and spherically symmetric case Padmanabhan2010 .
Moreover, we give a generalized definition for the internal energy of a black hole following the Hayward’s procedure Hayward1994b . Over a compact 2-surface $S$, Hayward introduces the Hamiltonian 2-form
@@
where $A_{0}$ is a dimensionless constant representing the area of a unit horizon [^12] We use $\sqrt{A/A_{0}}$ as it is the most natural unit of length., being $A_{0}=4\pi|g-1|$ in the non-flat case (where $g$ is the genus of the horizon) and $A_{0}=L_{x}L_{y}$ in the flat torus case [^13] This is the unique compact surface of constant Gaussian curvature equal to zero., where $L_{x}$ and $L_{y}$ are the length of the coordinate ranges after the identifications $x\sim x+L_{x}$ and $y\sim y+L_{y}$. In particular, introducing the Euler characteristic of the horizon, denoted by $\chi(H)$, Eq. (31) can be written as $U=\frac{\chi(H)}{2}\left(\frac{a}{2}\right)$ which coincides with the Padmanabhan’s definition of internal energy Padmanabhan2010 .
The quantity $\mbox{}^{\prime}\rho$ is constant over a stationary horizon Hayward1994d . Moreover, by direct computation $\Phi_{11}^{ph}+3\Lambda^{ph}$ depends only on the $r$ coordinate and the Gaussian curvature of the marginal surface, which we assumed to be constant. Thus, after using Eq. (31) in Eq. (28), this can be rewritten as
@@
As we commented before, Hayward’s definition for trapping gravity Hayward1994 , shown in Eq. (18) does not coincide with the surface gravity even in the context of static and spherically symmetric spacetimes , e.g. the Reissner-Nördstrom solution Fodor1996 ; Nielsen2008 . Our definition in Eq. (35) overcomes this problem and is valid for any static spacetime foliated by compact orientable 2-surfaces with constant Gaussian curvature.
@@
At this point, some comments about the thermodynamic volume of the black hole should be made. We first analyze the vacuum case in GR, where $\Phi_{11}^{ph}=0$, which admits a cosmological constant. Such a case has been widely studied, identifying the presence of a cosmological constant with a $P\mathrm{d}V$ term in the first law of black hole thermodynamics Kubizk2017 . Indeed, this allows the definition of the thermodynamic volume of the black hole, which may not coincide with the volume of a 3-manifold defined as having the horizon as its boundary (this 3-manifold may even be not well defined, for example in some hyperbolic cases). Thus, the thermodynamic volume should be defined as the conjugate thermodynamic variable to the pressure
@@
As commented in Sec. II, when a cosmological constant is considered, the mass of the black hole no longer coincides with its internal energy, but it is identified with the chemical enthalpy Kubizk2017 . This is,
@@
However, when returning to the general case with $\Phi_{11}\neq 0$, the resulting Penrose-Rindler Smarr formula is again the Smarr relation, but the Riemann decomposition is not the physical decomposition as the mass of the black hole is not $M=U+PV$ when $P$ is defined as in Eq. (36). An explicit example is shown below. Despite this, if the spacetime is stationary and asymptotically flat, the term $U+3PV$ coincides with the Komar energy of the spacetime Banerjee2010 .
@@
and noting that the left hand side term is equal to the black hole mass Villalba2020 this leads to
@@
For the vacuum case, the general action in $f(R)$ formalism is CapozzielloBEG
@@
Applying the variational principle, $\delta S=0$, the field equations for $f(R)$ gravity can be written as Eq. (8), where we have defined the effective stress-energy tensor as in Eq. (9).
@@
The first term coincides with the internal energy of the black hole in $f(R)$ gravity for the spherically symmetric case Sharif2012
@@
Additionally, the temperature remains the same independently of the theory (this is a expected result as the temperature is entirely determined by the geometry of the horizon) after identifying the corresponding entropy Faraoni2010
@@
V Towards the thermodynamics of more general configurations
Despite we introduced the Smarr formula within the GHP formalism for static spacetimes foliated by compact spacelike 2-surfaces of constant Gaussian curvature, the linking between Penrose and Rindler $K$-curvature and the Smarr formula suggests that the relation introduced in Eq. (22) is still related with the Smarr relation when integrated over a marginal surface for more general configurations. In this section, we briefly analyze [^15] We thank the anonymous referee for pointing this out. the encoding of the Smarr formula in the Penrose and Rindler $K$-curvature for some examples that are not contemplated in the initial spacetime geometric assumptions.
In particular, we study the charged Taub–NUT–AdS spacetime. The uncharged case is characterized by an extra parameter, $n$, which plays a role analogous to a gravomagnetic charge. In the metric there is a crossed term combining the time coordinate together with an angular one, reflecting a frame-dragging effect of the spacetime and the presence of so-called Misner strings singularities extending along the polar axes. For a detailed introduction see HawkingEllis .
The charged Taub–NUT–AdS spacetime is a solution of Einstein–Maxwell gravity that generalizes the Taub–NUT–AdS geometry by incorporating electric and magnetic charges Taub1951 ; Newman1963 . In such a case, the line element is given by
$ds^{2}=f(r)\left(dt+2n\cos\theta\,d\phi\right)^{2}-f(r)^{-1}dr^{2}-(r^{2}+n^{2})\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),$ (52)
with
$f(r)=\dfrac{r^{2}-2mr-n^{2}+4n^{2}g^{2}+e^{2}}{r^{2}+n^{2}}-\dfrac{3n^{4}-6n^{2}r^{2}-r^{4}}{l^{2}(r^{2}+n^{2})},$ (53)
where $e$ is the elerctric charge, $g$ is the magnetic charge and $l$ is the AdS radius.
In contrast with the ansatz introduced in Sec. IV, the principal null directions are rotating but again geodesic and shear-free, which implies $\omega\neq 0$ and $\kappa=\sigma=0$. Recalling that Eq. (22) is a general relation valid for any spacetime, which is simplified to
$k_{g}-2\tau\overline{\tau}-2\,\mathrm{Re}(\kappa\kappa^{\prime})-2\,\mathrm{Re}(\mbox{}^{\prime}\rho)+2\,\mathrm{Re}(\eth^{\prime}\tau)=2(\Phi_{11}^{ph}+3\Lambda^{ph}),$ (54)
when evaluated at the horizon, thus it leads to the expression
$k_{g}-2\,\mathrm{Re}(\mbox{}^{\prime}\rho)+2\,\mathrm{Re}(\eth^{\prime}\tau)=2(\Phi_{11}^{ph}+3\Lambda^{ph}),$ (55)
as $\tau=0$ for charged Taub-NUT AdS spacetimes. Hence, when integrating over the marginal surface we obtain
$2\pi\chi(H)-2\,\mathrm{Re}(\mbox{}^{\prime}\rho)A=2(\Phi_{11}^{ph}+3\Lambda^{ph})A,$ (56)
as $\mbox{}^{\prime}\rho,\Phi_{11}^{ph}$ and $\Lambda^{ph}$ are constant over the marginal surfaces and ð$\eth^{\prime}\tau$ is a gradient which has zero integral Hawking1972 .
Following a similar discussion that the one introduced in Sec. IV, Eq. (56) cannot be identified with the Smarr formula as it does not have energy units. Therefore, it is necessary to multiply the previous relation by a length dimension.
In contrast with the previous examples, the finding problem in this case is that Eq. (34) does not correctly defines the temperature of a charged Taub-NUT AdS black hole. This result suggests that the factor $\sqrt{A/A_{0}}$ is not valid for such a general case. Despite this, the Smarr formula can be recovered by multiplying Eq. (56) by $\frac{1}{8\pi}\frac{r_{+}^{2}+n^{2}}{r_{+}}$ (that is, we are replacing $\sqrt{A/A_{0}}\to\frac{r_{+}^{2}+n^{2}}{r_{+}}$), leading to
$\frac{\chi(H)}{4}\frac{r_{+}^{2}+n^{2}}{r_{+}}-\frac{1}{4\pi}\frac{r_{+}^{2}+n^{2}}{r_{+}}\mathrm{Re}(\mbox{}^{\prime}\rho)\,A=\frac{1}{4\pi}\frac{r_{+}^{2}+n^{2}}{r_{+}}(\Phi_{11}^{ph}+3\Lambda^{ph})A.$ (57)
The geometric motivation and interpretation of the factor $\frac{r_{+}^{2}+n^{2}}{r_{+}}$ is beyond the scope of this manuscript and it motivates an interesting path for linking Eq. (22) with the Smarr formula in even more generic configurations.
Note that $\chi(H)=2$ as the cross sections of the horizon are topological 2-spheres. Thus, the first term in Eq. (57) is given by $\frac{r_{+}^{2}+n^{2}}{2r_{+}}$, which coincides with the black hole mass for the case where no cosmological constant is considered Liu2024 . In such a case, as previously introduced in the manuscript, the mass is identified with the internal energy leading to
$U=\dfrac{r_{+}^{2}+n^{2}}{2r_{+}}.$ (58)
Moreover, the black hole temperature can be identified as
$T=\dfrac{\hbar}{2\pi}\dfrac{r_{+}^{2}+n^{2}}{r_{+}}\mathrm{Re}(\mbox{}^{\prime}\rho),$ (59)
after identifying the Bekenstein-Hawking entropy term $S=\frac{A}{4\hbar}$. In particular, a direct computation reveals that
$U=2TS+\dfrac{1}{4\pi}(\Phi_{11}^{ph}+3\Lambda^{ph})A,$ (60)
which implies that Eq. (57) is identically to
$M=2(TS+\psi N-PV)+\phi_{e}Q_{e}+\phi_{m}Q_{m},$ (61)
where all the well-known thermodynamic variables are described in Bordo2020 ; Hennigar2019 .
That is, we just showed that Eq. (22) is again related with the Smarr formula even in more general ansatz, although finding the necessary physical decomposition can be quite tedious for each particular case.
VI Final remarks and comments
@@
$k_{g}=K+\overline{K}\Leftrightarrow M=2TS+\Phi Q-2PV.$ (62)
@@
Even more, this framework allow us to study exotic topologies, obtaining some constraints on black hole existence in particular scenarios. For instance, the Penrose-Rindler Smarr formula implies that only spherical black holes with non-zero temperature can exist in General Relativity under the dominant energy condition. This recovers the well-known topology law for black holes Hayward1994b . Furthermore, the generalized definition of internal energy reveals that flat toroidal black holes have zero internal energy, while some hyperbolic black holes exhibit negative internal energy, which has been pointed out in previous works Villalba2020 .
The GHP formalism was initially developed for four-dimensional spacetimes, but it has been recently generalized to higher dimensions Durkee2010 . A key distinction in higher-dimensional analyses is the absence of a Gauss-Bonnet theorem, which plays a crucial role in relating black hole geometry to the Smarr law. Nevertheless, this extension suggests the possibility of analyzing thermodynamic formulations in higher-dimensional settings.
It is important to remark that the temperature we have proposed is not valid for the stationary axisymmetric case. In fact, despite the surface gravity over the Kerr black hole horizon must be constant due to the horizon topology, the quantity $\mbox{}^{\prime}\rho$ is not constant over the horizon (as it depends on angular coordinates), which suggests a revised generalization of trapping gravity for stationary spacetimes. A similar issue arises in the outlined charged Taub-NUT AdS case, where the presence of NUT charge further complicates the decomposition of relevant horizon quantities into the physical quantities. The study of this procedure applied to a more general ansatz is deferred to future work, including the connection between the Penrose and Rindler formula and the Smarr formula in Kerr-Newmann black holes or even in more generic configurations.
@@ Acknowledgments
We acknowledge financial support from the Generalitat Valenciana through PROMETEO PROJECT CIPROM/2022/13. A. G. acknowledges Fundación Humanismo y Ciencia for financial support. A.G. acknowledges Martuan for continuous support. We acknowledge the anonymous referees for their valuable comments and suggestions, which helped to improve this work.